OpenStudy (fearbigjjrob):

Question will be below, I think I have the answer, but am not 100% sure..

OpenStudy (fearbigjjrob):

When added, what is the simplified form of \[\left( 6+\sqrt{-4} \right)+\left( -15-\sqrt{-25} \right)\]? Write the answer in standard form, a + bi, where a and b are real numbers.

OpenStudy (fearbigjjrob):

The answer choices are A)-9-3i B)9-isqrt21 C)9+7i D)-9+7i

OpenStudy (mathmate):

@FearBigJJRob The best way is to post your answer for a check.

OpenStudy (fearbigjjrob):

@mhchen that is 1-21i ? I believe? the answer I believe it to be is either A or D, but am not sure how to finish the ending off

OpenStudy (fearbigjjrob):

I think it goes with the original equation becoming -9 I know for sure as the first, but what do I do with the negative square roots? My lesson is not telling me this..would I substitute a -2 in for sqrt-4 and a -5 for sqrt-25? @mhchen

OpenStudy (mathmate):

@FearBigJJRob I believe it is much faster to multiply it out (FOIL) and simplifying than to make guesses to the answers.

OpenStudy (fearbigjjrob):

Mathmate I did it how me lesson and the help video gave me, however they never showed us how to work with the -sqrt numbers... I think I have seen it somewhere before but I forgot, thats what I needed help on

OpenStudy (fearbigjjrob):

@mathmale do you know how to work this out? but for the equation\[\left( 7+\sqrt{-16} \right)=\left( -3-\sqrt{-36} \right)\] can you help me with this? either by showing me how to fully work this equation out or the same one with different numbers please? Im stuck.

OpenStudy (fearbigjjrob):

@whpalmer4 can you help me?

OpenStudy (mathmate):

\(\sqrt{-5}=\sqrt{-1\times 5}=\sqrt{-1}\times\sqrt{5}=\sqrt{5}i\) where \(i=\sqrt{-1}\) so \(\left( 6+\sqrt{-4} \right)+\left( -15-\sqrt{-25} \right)\) =\(\left( 6+\sqrt{4}i \right)+\left( -15-\sqrt{25}i \right)\) =\(\left( 6+2i \right)+\left( -15-5i \right)\) = ... I'll let you finish the sum.

OpenStudy (fearbigjjrob):

-9-3i thanks @mathmate now that I know sqrt of a negative number is just sqrt of that number plus i, thank you

OpenStudy (mathmate):

You're welcome! :)